OpenStudy (anonymous):

HELP!! A dump truck can haul 16 loads of dirt in 4 hours. How long will the dump truck spend hauling 25 loads? Options- (blank box)

OpenStudy (igreen):

16 loads/ 4 hours = 25 loads/ x hours x hours = (25 *4)/16 x hours = 100/16 = 6.25 hours, or 6 hours and 15 minutes Take a printscreen of your question though..

OpenStudy (anonymous):

Why? I have a BLANK box!!

OpenStudy (anonymous):

What is the final answer???

OpenStudy (igreen):

I know but just do it..

OpenStudy (igreen):

Take the printscreen, because I don't know if it wants you to put it in hours, minutes, or seconds..

OpenStudy (igreen):

Right, so the answer is 6.25.

OpenStudy (igreen):

The 'h' after the box tells me that it has to be in hours.

OpenStudy (anonymous):

ok